2 less than 3 times a number is the same as 3 more than 2 times the number. find the number.
Answer by ankor@dixie-net.com(22740)

Let x = "the number"
:
Write exactly what it says:
"2 less than 3 times a number is the same as 3 more than 2 times the number."
3x - 2 = 2x + 3
:
find the number.
3x - 2x - 3 + 2
x = 5
:
:
Check:
3(5) - 2 = 2(5) + 3